OpenStudy (anonymous):

|4x – 8| – 3x = 6x + 2

OpenStudy (anonymous):

|4x – 8| – 3x = 6x + 2

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

add a 3x to both sides to get \[|4x-8| = 9x + 2\] break the problem into 2 parts like so\[4x-8 = 9x + 2\] and \[4x - 8 = -9x - 2\] solve for x on both equations. lets start with the first one: \[4x-8 = 9x + 2\] subtract a 4x on both sides to get \[-8 = 5x + 2\] now subtract a 2 on both sides to get \[-10 = 5x\] divide by 5 on both sides, the first value of x = -2... now lets solve the second equation \[4x-8 = -9x -2\] add a 9x on both sides you get \[13x - 8 = -2\] then, add an 8 to both sides to get \[13x = 6\] divide by 13 on both side to get \[x = 6/13\]